$SPAD/src/input cclass.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool cclass.output
)set message test on
)set message auto off
)clear all
-- Input generated from CharacterClassXmpPage
--S 1 of 16
cl1 := charClass [char "a", char "e", char "i", char "o", char "u", char "y"]
--R
--R
--R   (1)  "aeiouy"
--R                                                         Type: CharacterClass
--E 1
--S 2 of 16
cl2 := charClass "bcdfghjklmnpqrstvwxyz"
--R
--R
--R   (2)  "bcdfghjklmnpqrstvwxyz"
--R                                                         Type: CharacterClass
--E 2
--S 3 of 16
digit()
--R
--R
--R   (3)  "0123456789"
--R                                                         Type: CharacterClass
--E 3
--S 4 of 16
hexDigit()
--R
--R
--R   (4)  "0123456789ABCDEFabcdef"
--R                                                         Type: CharacterClass
--E 4

--S 5 of 16
upperCase()
--R
--R
--R   (5)  "ABCDEFGHIJKLMNOPQRSTUVWXYZ"
--R                                                         Type: CharacterClass
--E 5
--S 6 of 16
lowerCase()
--R
--R
--R   (6)  "abcdefghijklmnopqrstuvwxyz"
--R                                                         Type: CharacterClass
--E 6
--S 7 of 16
alphabetic()
--R
--R
--R   (7)  "ABCDEFGHIJKLMNOPQRSTUVWXYZabcdefghijklmnopqrstuvwxyz"
--R                                                         Type: CharacterClass
--E 7
--S 8 of 16
alphanumeric()
--R
--R
--R   (8)  "0123456789ABCDEFGHIJKLMNOPQRSTUVWXYZabcdefghijklmnopqrstuvwxyz"
--R                                                         Type: CharacterClass
--E 8
--S 9 of 16
member?(char "a", cl1)
--R
--R
--R   (9)  true
--R                                                                Type: Boolean
--E 9
--S 10 of 16
member?(char "a", cl2)
--R
--R
--R   (10)  false
--R                                                                Type: Boolean
--E 10
--S 11 of 16
intersect(cl1, cl2)

--R
--R
--R   (11)  "y"
--R                                                         Type: CharacterClass
--E 11
--S 12 of 16
union(cl1,cl2)
--R
--R
--R   (12)  "abcdefghijklmnopqrstuvwxyz"
--R                                                         Type: CharacterClass
--E 12
--S 13 of 16
difference(cl1,cl2)
--R
--R
--R   (13)  "aeiou"
--R                                                         Type: CharacterClass
--E 13
--S 14 of 16
intersect(complement(cl1),cl2)
--R
--R
--R   (14)  "bcdfghjklmnpqrstvwxz"
--R                                                         Type: CharacterClass
--E 14
--S 15 of 16
insert!(char "a", cl2)
--R
--R
--R   (15)  "abcdfghjklmnpqrstvwxyz"
--R                                                         Type: CharacterClass
--E 15
--S 16 of 16
remove!(char "b", cl2)
--R
--R
--R   (16)  "acdfghjklmnpqrstvwxyz"
--R                                                         Type: CharacterClass
--E 16
)spool
)lisp (bye)

———-

References^{[1] nothing}

6